Shift Summary for Eve shift, 09/29/99

Sydney A. Taegar (taegar@physics.arizona.edu)
Wed, 29 Sep 1999 22:06:41 -0700 (MST)

Shift Summary for Eve shift, 09/29/99

On Shift: Wah and Sydney
Also present: BobT, JasonH, Valeri, RickK

Hours of beam delivered at all intensities: 1h
Hours of beam delivered at nominal intensity: 1h
Hours of E832 data (special runs not included): 0.7h
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

DDD1 veto due to sick ktev1 L3 processes.

========================================================================
Runs ended this shift:
Run 14757 - 25.5M events - BEAM_799

Tapes brought to FCC this shift:
UPA043-UPA048
UPB043-UPB048
UPC043-UPC048
UPS015-UPS016

========================================================================
Accesses during shift:

Cabibbo et al. on tour.
BobT, Valeri, Wah fix broken glue joint on V' counter 13.
RickT moves spill scope camera upstairs.

========================================================================
Overview of shift:

Beam was present for the first hour of shift but we lost
the last twenty minutes due to DDD1 veto. Ktev1 was processing
events at an *extremely* slow rate. We decided to stop the
run which subsequently hung, necessitating a fresh start.

JasonH changed HVs for CsI channels 228,266,1533,2691,3053,3134,3058
to fix high/low threshold and inefficient channels.

BobT, Valeri, and Wah fixed V' counter 13.

On consultation with Leo, BobT and JasonH set up fake spill signal to
allow various debugging tasks to continue in the absence of
the accelerator spill signal.

RickK improved RC10 src timing and, once again, fixed B16 by
plugging/unplugging cables. B16 is now back in COSMICS configuration.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Do PED run before beam.

Take LASER SCAN if no beam for > 1hr.